OpenStudy (anonymous):

find the distance between points A(7,3) and B(4,9) to the nearest tenth and express your answer in simplest form.

OpenStudy (anonymous):

\[AB=\sqrt{(9-3)^2+(7-4)^2}\]\[=\sqrt{36+9}=\sqrt{45}=3\sqrt{5}\]\to the nearest tenth\[\approx 6.7\]

OpenStudy (anonymous):

\[3\sqrt{5}\]is the exact simplified answer

OpenStudy (anonymous):

thxss